# Routoo: Learning to Route to Large Language Models Effectively

### v2

@@ 3. Increment budget by 1
Algorithm 1 Universe Construction.
The algorithm includes the following steps:
@@
    Performance-to-Cost Ratio: For each query $q_{j}$ and model $e_{i}$, calculate the performance-to-cost ratio. Let $\operatorname{\hat{eval}}(e_{i},q_{j})$ be the estimated evaluation score from the LLM and $c_{i}$ be the cost for model $e_{i}$. The ratio is calculated as:
    $\operatorname{ratio}_{ij}=\frac{\operatorname{\hat{eval}}(e_{i},q_{j})}{c_{i}^{\alpha}}$
@@
Figure 3: Pipeline of synthetic data generation of Routoo.
@@ 3.4 Universe Constructor
Given the vast array of LLMs available for text generation [^4] According to the Huggingface platform (https://huggingface.co), there are nearly 47,000 models available for the text generation task., and considering the practical constraints on the number of models that can be actively served, we propose an optimization approach to construct a complementary universe of models. This universe aims to maximize performance by ensuring that the selected models are complementary to each other.
The objective is to select a subset of LLMs that, collectively, provide the best coverage and performance across a set of queries. Formally, given a set of models $e_{1},e_{2},\ldots,e_{M}$, a set of queries $q_{1},q_{2},\ldots,q_{L}$, and the evaluations of these queries on the models $\operatorname{eval}(e_{i}(q_{j}))$, we seek to find the optimal subset of models that maximizes the average best response for the queries. The optimization problem is formulated as:
$\max_{U\subseteq M,|U|=k}S(U)=\frac{1}{L}\sum_{j=1}^{L}\max_{i\in U}\leavevmode\nobreak\ \operatorname{eval}(e_{i}(q_{j}))$ (4)
where $k$ is the desired number of serving models, $M$ is the set of all available LLMs, and $U$ is the subset of selected models from $M$. The function $S(U)$ represents the highest score achievable by the set $U$, quantifying the combined performance of the selected experts.
Given the submodular nature of the maximum operation in our objective function, a greedy algorithm (see Algorithm 1) can be employed to find an approximate solution with acceptable error bounds Krause and Golovin (2014). This method is particularly effective for large-scale problems where exact optimization is computationally prohibitive.
Further details on the generation of synthetic data for the universe construction are described in Section 4.2.
@@ 4.1 Experiment Setting
Evaluation Setting.
The evaluation of baselines and Routoo models is conducted using MMLU benchmark Hendrycks et al. (2021), which comprises multiple-choice questions across 57 diverse domains, such as mathematics, law, computer science, biology, and US history. To be compatible with OpenLLM Leaderboard Beeching et al. (2023), we use Eleuther AI Harness Gao et al. (2021) to evaluate our models.  [^5] Specifically, the following branch is used: https://github.com/EleutherAI/lm-evaluation-harness/tree/b281b0921b636bc36ad05c0b0b0763bd6dd43463. Precisely, it calculates the likelihood of each choice in the question, and selects the answer with maximum likelihood, then ’accuracy’ is used as the evaluation metric. The overall performance is calculated as the average accuracy of the model in 57 domains.
Baselines.
Our comparison includes a range of both open-source and closed-source LLMs. These comprise LLaMa2 Touvron et al. (2023) models with 7b, 13b, and 70b parameters, Mistral 7b Jiang et al. (2023a), Mixtral 8x7b Jiang et al. (2024) (employing token-level MoE), alongside with GPT3.5 and GPT4 OpenAI et al. (2024) as closed-source models.
Architecture Setting.
For the universe constructor, we use all available models of OpenLLM leaderboard Beeching et al. (2023)  [^6] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard. as the underlying LLMs. At the time of writing this paper, there are nearly 2,200 LLMs available on the leaderboard. We set the size of the universe as 56. For the performance predictor, we use Mistral 7b (v0.1) Jiang et al. (2023a)  [^7] https://huggingface.co/mistralai/Mistral-7B-v0.1. as the query encoder. The number of levels in the evaluation score is 2 for MMLU benchmark.  [^8] As accuracy metric is used for MMLU benchmark in OpenLLM leaderboard Beeching et al. (2023). For fine-tuning, we apply LoRA method Hu et al. (2021) with $r=1024,\leavevmode\nobreak\ \alpha=16,\leavevmode\nobreak\ dropout=0.05$ on query, key, and value matrices.
4.2 Training Data Preparation
The data generation pipeline for Routoo is depicted in Figure 3, featuring a query generator, a response generator, and an evaluation system (specifically the MMLU evaluation system). The process begins with the specialized question generator; questions were originated by two approaches: Filtering available open-source datasets and generating queries using GPT4 model OpenAI et al. (2024).
Filtering Available Datasets
: We began by collecting various multiple-choice QA datasets, such as ARC Clark et al. (2018), MC-TEST Richardson et al. (2013), OBQA Mihaylov et al. (2018), RACE Lai et al. (2017), and TruthfulQA Lin et al. (2022). To identify questions with a high training signal (i.e., difficulty), we employed SOLAR-10.7B-v1.0 Kim et al. (2024)  [^9] Available in Huggingface platform: https://huggingface.co/upstage/SOLAR-10.7B-v1.0. We chose this model, as it is performing relatively well on MMLU section of OpenLLM benchmark Beeching et al. (2023). on this aggregated dataset, and calculate the evaluation score. Questions where the model performed poorly were retained, yielding 45,645 challenging training samples from an initial pool of 101,434. Additionally, 10,000 simpler questions were randomly selected from the initial dataset, bringing the total to 55,645 queries.
Generation with GPT4
: To diversify our training dataset further, we generated 20,000 synthetic queries using the GPT4 model. These queries were based on domain and seed data consisting of five random questions from our dataset mentioned above. Detailed specifications of the input prompts used for generation are documented in Appendix B.
After generating the queries, they are passed to the response generator to generate outputs from LLMs in the universe set. Finally, for a given query $q_{j}$ and the response from model $e_{i}$, the evaluation system computes $\operatorname{eval}(e_{i}(q_{j}))$.
The development set of Universe Constructor is created by running a response generator and evaluation system on available LLMs from the OpenLLM benchmark Beeching et al. (2023) (approximately 2,200) for 1,000 randomly selected queries from our training dataset.
@@ 4.3 Main Results
We present two variations of Routoo model: Routoo (open-source) and Routoo (mix). The former refers to a model that leverages 7B, 13B ,and 34B open-source models available on Huggingface as the input to our universe constructor. By additionally integrating GPT4 OpenAI et al. (2024) to underlying models of Routoo (open-source), we create Routoo (mix).
@@
Comparison of our models and baselines are illustrated in Table 1.101010Inference costs are calculated based on price documentation of the following providers at the time of writing the paper: https://www.together.ai, https://openai.com. For GPT3.5 and GPT4 costs, the average of input and output costs are considered. The cost of Routoo (open-source) is adjusted to match Mixtral 8x7b Jiang et al. (2023a), the best generic open-source LLM (at the time of writing this paper). Our model significantly outperforms Mixtral by 5.27% absolute points with the same inference cost. Also, our underlying models can be executed on 1 GPU (e.g. A100 with 40 GB memory), while Mixtral model requires access to high-end computing resources  (e.g. GPUs with 100 GB memory). Compared to LLaMa2 70B, our model achieves significantly better performance (+6%), while reducing the cost by 33%. Impressively, Routoo (open-source) achieves competitive performance with GPT3.5 while reducing the cost by 73.3%.
Compared to closed-source LLMs, our Routoo (mix) model significantly outperforms GPT3.5 by 14.9% absolute point. Routoo (mix) reaches competitive performance with GPT4, while reducing the cost by almost 50%.
Figure 6: Routing distributions of different variants of Routoo models.
Cost-Aware Decoding.
Given that the cost-aware decoding module can compute different routing distributions within a specified budget, we illustrate the curve of performance-cost in Figure 4 for both open-source and closed-source baselines, alongside with variations of Routoo models. Notably, our Routoo (open-source) reaches the performance of Mixtral 8x7b Jiang et al. (2024) while decreasing the cost by 33%. Interestingly, Routoo (mix) achieves better performance than GPT4 (86.64 vs. 86.40) while reducing the inference cost by nearly 25%.
In summary, as our cost-aware decoder effectively offers an optimal trade-off between cost and performance, both the open-source and mix variants of the Routoo model provide a better performance (or cost) compared to using individual LLMs for queries during inference.
@@
To further investigate the source of improvement in our models, we illustrates the distributions of performances on 17 sub-categories Hendrycks et al. (2021) in Figure 5. A standout area of success is in STEM domains, such as mathematics and computer science, where the Routoo particularly excels. This impressive performance is largely attributed to the incorporation of specialized small models (around 7b parameters) that are fine-tuned for tasks in mathematics Yu et al. (2024); Shao et al. (2024) and coding Rozière et al. (2024); Guo et al. (2024) by the community. Furthermore, this approach facilitates the identification of domains where there is a scarcity of experts. Then, future research can focus on improving the performance of these areas by developing domain-specific effective experts.
@@
Figure 6 presents the aggregated percentage usage of expert models by size for each Routoo pricing tier. It reveals that higher-priced options tend to utilize larger models more frequently. A substantial inclusion of effective smaller, 7-billion parameter expert models significantly enhances the cost-to-performance efficiency. This suggests that strategically increasing the use of such smaller experts could offer a more economical solution while maintaining high-quality outputs.
Finally, the cost of training a router is significantly lower than building foundational models e.g. GPT4 OpenAI et al. (2024) and Mixtral Jiang et al. (2024). This could pave a new path for building new frontiers at a much lower cost by integrating knowledge of current LLMs.
@@ 5 Conclusion and Future Work
In this paper, we proposed our Routoo architecture, a lightweight LLM-based model that is designed to inteligently routes the input query to the most suitable expert model given other constraints e.g. cost. We evaluated our architecture on MMLU benchmark Hendrycks et al. (2021), which is a MCQA dataset with 57 different domains coverage. Routoo (open-source) achieved competitive performance with Mixtral 8x7b model Jiang et al. (2024) with two-thirds of the inference cost. Increasing the budget limitation allows Routoo (open-source) to outperform Mixtral model by 5% with the same level of cost. By integrating GPT4 OpenAI et al. (2024) model to underlying experts, our Routoo (mix) achieved competitive accuracy with GPT4 while reducing the inference cost by 50%, and even surpassing it while saving 25% of the cost. In general, our Routoo models provide an efficient trade-off between cost and performance during the inference.
In future, the Routoo’s ability to assess and understand the performance of existing models allows researchers to identify gaps in the AI landscape. It pinpoints domains where no existing expert excels or where larger models are inefficiently juggling tasks. This insight is invaluable, enabling us to strategically develop domain-specific models where they are needed most.
Limitations
Due to GPU limitations, we fine-tuned our Routoo models using the LoRA technique Hu et al. (2021). For better performance, full fine-tuning methods could be employed to train our Routoo models. Our cost-aware decoder used a greedy technique to account for inference costs, though more comprehensive methods could better model the relationship between performance and cost. Resource constraints prevented us from using LLMs with 70 billion parameters as our underlying models. Incorporating these larger models, especially in the Routoo (mix) model, yields better performance compared to GPT4. Additionally, integrating other closed-source LLMs e.g. Claude  [^11] https://www.anthropic.com/news/claude-3-family and Gemini  [^12] https://deepmind.google/technologies/gemini/ variants into our Routoo models would increase the diversity of base LLMs.
@@ Appendix A Optimal Routing of Open-source LLMs on MMLU
In this experiment, we utilize all publicly available LLMs from the OpenLLM benchmark Beeching et al. (2023) that have fewer than 34 billion parameters. Given that each query is directed to the most effective and cost-efficient model by an ideal performance predictor module within Routoo, the upper bound for the Routoo model in this scenario is nearly 97.5%. The following table shows the distribution of usage among LLMs of varying sizes:
@@
Table 2: The Distribution of usage among underlying LLMs with different sizes on MMLU Hendrycks et al. (2021) test set.

### v3

@@ 3. Increment budget by 1
Algorithm 1 Greedy algorithm for the optimization of universe construction.
where $\operatorname{Pred}(m_{i},q_{j})$ is the predictive model. Similar to $s_{m_{i},q_{j}}$, $\hat{s}_{m_{i},q_{j}}$ is an integer that ranges from 0 to $K-1$, 0 indicates an unacceptable result and $K-1$ represents the optimal outcome.
The process of $\operatorname{Pred}(m_{i},q_{j})$ is as follows:
$\begin{split}\begin{cases}h_{q_{j}}=\operatorname{Enc}(q_{j})\\ h_{m_{i}}=\operatorname{Emb}(m_{i})\\ \hat{s}_{m_{i},q_{j}}=\operatorname{Linear}(h_{q_{j}}-h_{m_{i}})\end{cases}\end{split}$
where $\operatorname{Enc}(.)$ is the encoder of the input query. Inspired by Radford et al. (2019), we use a decoder-only LLM as the query encoder and extract the embedding of the last token as the representation of the input query ($h_{q_{j}}\in\mathbb{R}^{h}$). $\operatorname{Emb}(.)$ is an embedding layer, where each embedding is assigned to a specific model. As a result, $h_{e_{i}}\in\mathbb{R}^{h}$ is the embedding representation of model $e_{i}$. $h$ is the hidden representation of the decoder-only model used. $\operatorname{Linear}(.)$ function is a projection matrix ($\mathbb{R}^{h\times K}$), computing the estimation score $\hat{s}_{m_{i},q_{j}}$.
The model is trained by minimising the cross-entropy loss function (Good, 1952) of $s_{m_{i},q_{j}}$ and $\hat{s}_{m_{i},q_{j}}$ over all $M$ models, and $N$ queries as:
$\operatorname{min}\frac{1}{N\times M}\sum_{i=1}^{M}\sum_{j=1}^{N}\operatorname{CE}(s_{m_{i},q_{j}},\hat{s}_{m_{i},q_{j}})$ (1)
where $\operatorname{CE}(.)$ is the cross-entropy loss.
During the inference time, $\operatorname{argmax}(.)$ function is applied for a given model and query.
3.3 Cost-Aware Selector
The second phase of our architecture involves the selection step, where the estimated scores from the performance predictor are used to determine the optimal assignment of models to queries. Our objective is to maximize the overall effectiveness of the responses within a given budget constraint. The optimization problem is formulated as follows:
$\max_{\pi}$ $\ \sum_{j=1}^{N}\hat{s}_{\pi(q_{j}),q_{j}}$
s.t. $\ \sum_{j=1}^{N}c_{\pi(q_{j}),q_{j}}\leq B$
where $\pi(.)$ determines the model assignments and $B$ represents the predefined budget constraint.
We propose a greedy algorithm to approximate this optimization. Noting the challenges of accurately estimating the exact cost of a model for a specific query, instead, we consider the average cost $c_{i}$ for a model $m_{i}$ responding to an average length query and response. The algorithm includes the following steps:
@@ 3.3 Cost-Aware Selector
    Performance-to-Cost Ratio: For each query $q_{j}$ and model $m_{i}$, calculate the performance-to-cost ratio. Let $\hat{s}_{m_{i},q_{j}}$ be the estimated correctness score of model $m_{i}$ for query $q_{j}$, and $c_{i}$ be the cost for model $m_{i}$. The ratio is calculated as:
    $r_{m_{i},q_{j}}=\frac{\hat{s}_{m_{i},q_{j}}}{c_{i}^{\alpha}}$
@@ 3.4 Universe Constructor
With the abundance of models and the practical limitations on hosting several models, we introduce an optimization approach to build a complementary subset of models ($m_{1},m_{2},\ldots,m_{M}$). The goal is to select models that are complementary to achieve the highest performance, given a predefined limitation on number of serving models.
The objective is to select a subset of models that collectively provide the best coverage and performance across a set of queries. Formally, given a big pool of models $\Omega:(m_{1},m_{2},\ldots,m_{\hat{M}})$, a set of queries $Q:(q_{1},q_{2},\ldots,q_{L})$, and correctness scores of the models for these queries $s_{m_{i},q_{j}}$, we seek to find the optimal subset of models, that maximizes the following optimization problem:
$\max_{U\subseteq\Omega,|U|=M}S(U)=\frac{1}{L}\sum_{j=1}^{L}\max_{i\in U}~{}s_{m_{i},q_{j}}$ (2)
where $M$ is the desired number of serving models, and $U$ is a subset of selected models from $\Omega$. The function $S(U)$ represents the highest score achievable by the set $U$, quantifying the combined performance of the selected models. The optimal subset $S(U^{*})$ is used as the set of models ($m_{1},m_{2},\ldots,m_{M})$ in the optimization of Section 3.1.
Given the submodular nature of the maximum operation in our objective function (Equation 2), a greedy algorithm (as illustrated in Algorithm 1) can be employed to find an approximate solution with acceptable error bounds (Krause & Golovin, 2014). This method is particularly effective for large-scale problems where exact optimization is computationally prohibitive.
@@ 4.1 Experiment Setting
Evaluation Setting.
The evaluation of baselines and Routoo models is conducted using MMLU benchmark (Hendrycks et al., 2021), which comprises multiple-choice questions across 57 diverse domains, such as mathematics, law, computer science, biology, and US history. To be compatible with OpenLLM Leaderboard (Beeching et al., 2023), we use Eleuther AI Harness (Gao et al., 2021) to evaluate our models.  [^5] Specifically, the following branch is used: https://github.com/EleutherAI/lm-evaluation-harness/tree/b281b0921b636bc36ad05c0b0b0763bd6dd43463. Precisely, it calculates the likelihood of each choice in the question, and selects the answer with maximum likelihood, then ’accuracy’ is used as the evaluation metric. The overall performance is calculated as the average accuracy of the model in 57 domains. Routoo models will be made publicly available for the AI community to test on various benchmarks.
Baselines.
Our comparison includes a range of both open-source and closed-source LLMs. These comprise LLaMa2 (Touvron et al., 2023) models with 7b, 13b, and 70b parameters, Mistral 7b (Jiang et al., 2023a), Mixtral 8x7b (Jiang et al., 2024) (employing token-level MoE), alongside with GPT3.5 and GPT4 (OpenAI et al., 2024) as closed-source models. We could not find any publicly available trained models for routing or model integration. To the best of our knowledge, there are currently no open-source models that perform routing or dynamic selection among multiple LLMs that we could directly compare against.
Architecture Setting.
For the pool of universe constructor ($\Omega$), we use top 1,000 available models of OpenLLM leaderboard (Beeching et al., 2023)  [^6] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard., Then, we set the maximum number of serving models ($M$) to 56 in Equation 2, meaning that we use 56 models for the routing optimization, defined in Section 3.1. For the performance predictor, we use Mistral 7b (v0.1) (Jiang et al., 2023a)  [^7] https://huggingface.co/mistralai/Mistral-7B-v0.1. as the query encoder. The number of levels in the evaluation score ($K$) is set to 2 for MMLU benchmark, meaning that $s_{m_{i},q_{j}}$ and $\hat{s}_{m_{i},q_{j}}$ are either 0 or 1.  [^8] As accuracy metric is used for MMLU benchmark in OpenLLM leaderboard (Beeching et al., 2023). For fine-tuning, we apply LoRA method (Hu et al., 2021) with $r=1024,~{}\alpha=16,~{}\text{dropout}=0.05$ on query, key, and value matrices.
4.2 Training Data Preparation
Since the MMLU dataset (Hendrycks et al., 2021) lacks a training set, we build synthetic data to train the models. The synthetic questions are originated by two approaches: Filtering available open-source datasets and generating synthetic queries using GPT4 model (OpenAI et al., 2024).
Filtering Available Datasets.
We begin by collecting various multiple-choice QA datasets, such as ARC (Clark et al., 2018), MC-TEST (Richardson et al., 2013), OBQA (Mihaylov et al., 2018), RACE (Lai et al., 2017), and TruthfulQA (Lin et al., 2022). To identify questions with a high training signal (i.e., difficulty), we employed SOLAR-10.7B-v1.0 (Kim et al., 2024)  [^9] Available in Huggingface platform: https://huggingface.co/upstage/SOLAR-10.7B-v1.0. We chose this model, as it is performing relatively well on MMLU section of OpenLLM benchmark (Beeching et al., 2023). on this aggregated dataset, and calculate the evaluation score. [^10] MMLU evaluation system (introduced in Section 4.1) is used. Queries on which the model under-performed (accuracy of 0) were retained, resulting in a set of 45,645 challenging training samples from an initial pool of 101,434. Additionally, 10,000 simpler questions were randomly selected from the initial dataset, bringing the total to 55,645 queries.
@@ 4.2 Training Data Preparation
Generation with GPT4.
To diversify our training dataset further, we generated 20,000 synthetic queries using GPT4 model. These queries are generated by using seed data, which randomly sampled from the dataset mentioned above. Detailed specifications of the input prompts used for generation are documented in Appendix B.
Final set of queries contains nearly 75,000 samples. Datatrove library [^11] https://github.com/huggingface/datatrove. is used for decontamination. The set of queries in Universe Constructor ($Q:(q_{1},q_{2},\ldots,q_{L})$) is created by randomly sampling 1,000 queries from the training dataset.
@@ 4.3 Main Results
We present two variations of Routoo model: Routoo (open-source) and Routoo (mix). The former refers to a model that leverages 7b, 13b ,and 34b open-source models available on Huggingface as the input to our universe constructor. By additionally integrating GPT4 (OpenAI et al., 2024) to underlying models of Routoo (open-source), we create Routoo (mix).
@@
Comparison of our models and baselines are illustrated in Table 1.121212Inference costs are calculated based on price documentation of the following providers at the time of writing the paper: https://www.together.ai, https://openai.com. For GPT3.5 and GPT4 costs, the average of input and output costs are considered. The cost of Routoo (open-source) is adjusted to match Mixtral 8x7b (Jiang et al., 2023a), the best generic open-source LLM (at the time of writing this paper). Our model significantly outperforms Mixtral 8x7b by 5.27% absolute points with the same inference cost. Also, our underlying models can be executed on 1 GPU (e.g. A100 with 40 GB memory), while Mixtral model requires access to high-end computing resources  (e.g. GPUs with 100 GB memory). Compared to LLaMa2 70b, our model achieves significantly better performance (+6%), while reducing the cost by 33%. Impressively, Routoo (open-source) achieves competitive performance with GPT3.5 while reducing the cost by 73.3%. Compared to closed-source LLMs, our Routoo (mix) model significantly outperforms GPT3.5 by 14.9% absolute point. Routoo (mix) reaches competitive performance with GPT4, while reducing the cost by almost 50%.
Figure 4: Routing distributions of different variants of Routoo models. The inference cost is provided for 1 million tokens.
Cost-Aware Selection.
[…]
In summary, our cost-aware selector effectively offers an optimal trade-off between cost and performance, enabling both open-source and mix variants of the Routoo model to outperform individual LLMs in terms of performance (or cost) during inference.
@@
To further investigate the source of improvement in our models, we illustrates the distributions of performances on 17 sub-categories (Hendrycks et al., 2021) in Figure 3. A standout area of success is in STEM domains, such as mathematics and computer science, where Routoo particularly excels. This impressive performance is largely attributed to the incorporation of specialized small models (around 7b parameters) that are fine-tuned for tasks in mathematics (Yu et al., 2024; Shao et al., 2024) and coding (Rozière et al., 2024; Guo et al., 2024) by the community. Furthermore, this approach facilitates the identification of domains where there is a scarcity of experts. Then, future research can focus on improving the performance of these areas by developing domain-specific effective experts.
@@
Figure 4 presents the aggregated percentage usage of expert models by size for each Routoo pricing tier (\$ per 1M tokens). It reveals that higher-priced options tend to utilize larger models more frequently. A substantial inclusion of effective smaller, 7-billion parameter expert models significantly enhances the cost-to-performance efficiency. This suggests that strategically increasing the use of such smaller experts could offer a more economical solution while maintaining high-quality outputs.
Finally, the cost of training a router is significantly lower than building foundational models e.g. GPT4 (OpenAI et al., 2024) and Mixtral (Jiang et al., 2024). This could pave a new path for building new frontiers at a much lower cost by integrating knowledge of current LLMs.
@@ 5 Conclusion and Future Work
In this paper, we proposed our Routoo architecture, a lightweight LLM-based model that is designed to inteligently routes the input query to the most suitable expert model given other constraints e.g. cost. We evaluated our architecture on MMLU benchmark (Hendrycks et al., 2021), which is a MCQA dataset with 57 different domains coverage. Routoo (open-source) achieved competitive performance with Mixtral 8x7b model (Jiang et al., 2024) with two-thirds of the inference cost. Increasing the budget limitation allows Routoo (open-source) to outperform Mixtral model by 5% with the same level of cost. By integrating GPT4 (OpenAI et al., 2024) model to underlying LLMs, our Routoo (mix) achieved competitive accuracy with GPT4 while reducing the inference cost by 50%, and even surpassing it while saving 25% of the cost. In general, our Routoo models provide an efficient trade-off between cost and performance during the inference.
In future, the Routoo’s ability to assess and understand the performance of existing models allows researchers to identify gaps in the AI landscape. It pinpoints domains where no existing expert excels or where larger models are inefficiently juggling tasks. This insight is invaluable, enabling us to strategically develop domain-specific models where they are needed most. Furthermore, our architecture facilitates easy integration of additional optimization criteria, including cost, speed, and privacy considerations.
@@ Appendix A Optimal Routing of Open-source LLMs on MMLU
In this experiment, we utilize all publicly available LLMs from the OpenLLM benchmark (Beeching et al., 2023) that have fewer than 34 billion parameters. Given that each query is directed to the most effective and cost-efficient model by an ideal performance predictor module within Routoo, the upper bound for the Routoo model in this scenario is nearly 97.5%. The following table shows the distribution of usage among LLMs of varying sizes:
@@
Table 2: The Distribution of usage among underlying LLMs with different sizes on MMLU (Hendrycks et al., 2021) test set.